Your company is considering the replacement of an old delivery van with a new one that is more efficient. The old van cost $40,000 when it was purchased 5 years ago. The old van is being depreciated using the simplified straight -line method over a useful life of 8 years. The old van could be sold today for $7000. The new van has an invoice price of $80,000 and it will cost $6000 to modify the van to carry the company's products. Cost savings from the use of the new van are expected to be $28,000 per year for 5 years, at which time the van will be sold for its estimated salvage value of $18,000. The new van will be depreciated using the simplified straight-line method over its 5-year useful life. The company's tax rate is 35%. Working capital is expected to increase by $5,000 at the inception of the project but this amount will be recaptured at the end of year five. What is the tax effect of selling the old machine?

a. a savings of $2,800

b. a savings of $2,450

c. additional taxes paid of $2,450

d. a tax savings of $1,400
Answer #1
Sale price of old van 7000
Book value = 40000-40000*5/8 = 15000
Loss on sale = 15000-7000 = 8000
Tax shield on loss = 8000*35% = 2800
Answer: Option [a] A savings of $2800